# Diffusion approximation of the stationary distribution of a two-level single server queue

### v4

@@ 4.2 Preliminary lemmas
We will prove Theorem 3.1 using the moment generating functions $\varphi^{(r)}_{i}(\theta_{2})$ and $\varphi^{(r)}_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$. Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}],$
@@
$g^{(r)}_{i,\theta}(x_{1},x_{2})=e^{-\eta^{(r)}_{1}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r)},\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
@@
$e^{\theta}\mathbb{E}\left[e^{-\eta^{(r)}_{i}(\theta)(T^{(r)}_{e,i}\wedge 1/r)}\right]=1,\quad i=1,2,\quad e^{-\theta}\mathbb{E}\left[e^{-\zeta^{(r)}(\theta)(T^{(r)}_{s}\wedge 1/r)}\right]=1.$ (4.8)
Note that $\eta^{(r)}_{i}(\theta)$ is the value of the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ at $e^{-\theta}$, so it is infinitely many differentiable function of $\theta$. Obviously, $\zeta^{(r)}(\theta)$ has the same property. Hence, twice differentiating the equations in (4.8), we can compute the Taylor expansions of them around the origin, which yields the following lemma. It is formally obtained in [3, Lemma 5.8].
@@
because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations bellow for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
Based on these results, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{{\boldsymbol{\theta}}}$. Then,
@@
Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]\right)$. Then, by (4.2) and (4.2),
@@
and, for $k=1,2$,
$\mathbb{E}[(\widehat{T}^{(r)}_{u})^{k}]=\mathbb{E}[(T^{(r)}_{u})^{k}]+o(r^{3-k}),\qquad u=(e,1),(e,2),s,$ (4.24)
$\mathbb{E}[(\widehat{R}^{(r)}_{v})^{k}]=\mathbb{E}[(R^{(r)}_{v})^{k}]+o(r^{2-k}),\qquad v=e,s,$ (4.25)
@@
In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly integrable.
@@
For $k=1$, multiply both sides of (4.5) and (4.5) by $\mu^{(r)}$ and $\lambda^{(r)}_{1}$, respectively, and take their difference. Since $c^{(r)}_{1}\mu^{(r)}-\lambda^{(r)}_{1}=r\mu^{(r)}b_{1}+o(r)$ by the assumption (3.c), $\mathbb{P}(R^{(r)}_{e}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.22) and $\mathbb{P}(R^{(r)}_{s}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.23), we have
@@
because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1. This proves (4.33). Similarly, (4.34) is obtained from (4.5) and (4.5), ∎
@@
(4.45) is immediate from Lemma 4.6 by applying (4.43) and (4.44). As for (4.46), we only prove its first equation because its second one is similarly proved. By the Schwartz’s inequality,
@@
$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad\leq\mathbb{E}[(T^{(r)}_{s})^{3}]+3c_{1}\mathbb{E}[(R^{(r)}_{s})^{2}1(R^{(r)}_{s}\geq 1/r,L^{(r)}\geq 1)]/\alpha^{(r)}$
$\quad=\mathbb{E}[(T^{(r)}_{s})^{3}]+\frac{c_{1}}{\min_{i=1,2}c_{i}}\mathbb{E}[(T^{(r)}_{s})^{3}1(T^{(r)}_{s}\geq 1/r)],$
which is uniformly bounded for $r\in(0.1]$, and therefore (4.3) and (4.44) prove (4.46). ∎
@@
Hence, (4.66) holds. Applying it to (4.61) and (4.64) for $i=1$, we obtain
@@
We next prove (ii). For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumption in (ii). First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$. In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{e}$ under $\mathbb{P}^{(r)}_{s}$, so, by Lemma 4.4 for $k=1$,
$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)$ $=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
@@
Since $\mu c_{i}=\lambda_{i}$, this can be written as
@@
because $\lim_{r\downarrow 0}\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=0$ and (4.44). On the other hand, if $b_{1}=0$, then (4.54) and (4.60) yield
@@
Since $c_{1}=c_{2}$, this proves (4.71). Similarly, if $b_{1}\not=0$, then (4.54) and (4.63) yield
@@
Again by $c_{1}=c_{2}$, this proves (4.72).
@@ 5 Concluding remarks
In this section, we intuitively answer two questions, what is a process limit for the 2-level $GI/G/1$ queue in heavy traffic and how to verify (3.11) in Remark 3.1.
@@ 5.1 Process limit
So far, in the previous versions of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered the stationary distribution using the generalized Ito formula for a convex function. We got a stationary equation similar to (5.1) for the $Z(\cdot)$, but could not completely obtained the stationary distribution although we had partial information which is compatible with Theorem 3.1. So, it is conjectured in the previous versions that
@@
During the reviewing process of this paper, the author have known that (a) is resolved by Atar et al. [1, Lemma 4.3] (see also [12] for a different proof), and (c) is affirmatively answered in [12]. Furthermore, the stationary distribution of $Z(t)$ which is obtained in [12] does not need any of the extra conditions in Theorem 3.1. This is not surprising because limiting operation for $r\downarrow 0$ is not needed to derive the stationary distribution of $Z(t)$. Thus, we need not further discuss about the conjectures (a) and (c). As for the conjecture (b), it is still open (see [12]).

### v5

@@ 4.2 Preliminary lemmas
We will prove Theorem 3.1 using the moment generating functions $\varphi^{(r)}_{i}(\theta_{i})$ and $\varphi^{(r)}_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$. Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{0})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(L^{(r)}>\ell^{(r)}_{0})],$
@@
$g^{(r)}_{i,\theta}(x_{1},x_{2})=e^{-\eta^{(r)}_{i}}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r),\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
@@
$e^{\theta}\mathbb{E}\left[e^{-\eta^{(r)}_{i}(\theta)(T^{(r)}_{e,i}\wedge 1/r)}\right]=1,\quad i=1,2,\quad e^{-\theta}\mathbb{E}\left[e^{-\zeta^{(r)}(\theta)(T^{(r)}_{s}\wedge 1/r)}\right]=1,$ (4.8)
because $\eta^{(r)}_{i}(\theta)$ is the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ taking the value $e^{-\theta}$ and $\zeta^{(r)}(\theta)$ is similarly determined. Furthermore, $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are infinitely many differentiable functions of $\theta$. Hence, twice differentiating the equations in (4.8), the following lemma can be obtained by their Taylor expansions around the origin. It is formally proved in [3, Lemma 5.8].
@@
because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
Based on these results, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{r{\boldsymbol{\theta}}}$. Then,
@@
Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]\right)$. Then, by (4.2) and (4.2),
@@
and, for $i=1,2$,
$|\mathbb{E}[(\widehat{T}^{(r)}_{u})^{k}]-\mathbb{E}[(T^{(r)}_{u})^{k}]|=\mathbb{E}[(T^{(r)}_{u})^{k}1(T^{(r)}_{u}>1/r)]=o(r^{3-k}),\quad u=(e,i),s,$ (4.24)
$|\mathbb{E}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}[(R^{(r)}_{e})^{k}]|=\mathbb{E}[(R^{(r)}_{v})^{k}1(R^{(r)}_{v}>1/r)]=o(r^{2-k}),\qquad v=e,s,$ (4.25)
@@
In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly integrable for $k=1,2$.
@@
For $k=1$, multiply both sides of (4.5) and (4.5) by $\mu^{(r)}$ and $\lambda^{(r)}_{1}$, respectively, and take their difference, then, noting that $c^{(r)}_{1}\mu^{(r)}-\lambda^{(r)}_{1}=r\mu^{(r)}b_{1}+o(r)$, we have
@@
because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.25). This proves (4.33). Similarly, (4.34) is obtained from (4.5) and (4.5), ∎
@@
The first equation of (4.45) is immediate from (4.2) and Lemma 4.6 together with (4.43) and (4.44), and the second equation follows from the first one by Lemma 4.1. As for (4.46), we only prove its first equation because its second one is similarly proved. By the Schwartz’s inequality,
@@
$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\leq\mathbb{E}[(T^{(r)}_{s})^{3}],$
which is uniformly bounded for $r\in(0,1]$, and therefore (4.3) and (4.44) prove (4.46). ∎
@@
Hence, (4.66) holds. Applying it to (4.61) and (4.64), we obtain
@@
We next prove (ii). For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumption in (ii). First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$. In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$, so, by Lemma 4.4 for $k=1$,
$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)\;=\ell^{(r)}_{0})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
@@
Since $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$ and $\mu c_{i}=\lambda_{i}$, this can be written as
@@
because $\lim_{r\downarrow 0}\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=0$ and (4.44). If $b_{1}=0$, then (4.54) and (4.60) yield
@@
Since $c_{1}=c_{2}$, this proves (4.71). Similarly, if $b_{1}\not=0$, then (4.54), (4.63) and (4.3) yield
@@
Again by $c_{1}=c_{2}$ and $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$, this proves (4.72).
@@ 5 Concluding remarks
In this section, we intuitively discuss two questions, what is a process limit for the 2-level $GI/G/1$ queue in heavy traffic and how to verify (3.12) in Remark 3.1.
@@ 5.1 Process limit
So far, in the original submission of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered the stationary distribution using the generalized Ito formula for a convex function. We got a stationary equation similar to (5.1) for the $Z(\cdot)$, but could not completely obtained the stationary distribution although we had partial information which is compatible with Theorem 3.1. So, it was conjectured that
@@
During the reviewing process of this paper, the author have known that (a) is resolved by Atar et al. [1, Lemma 4.3] (see also [12] for a different proof), while the author affirmatively answers (c) in [12]. Those results are obtained for more general reflecting diffusion processes than $Z(\cdot)$, which are hard to include in this paper. Thus, we refer to [12] for them, and will not further discuss about the conjectures (a) and (c). As for the conjecture (b), it is still open (see [12]).